At a depth in water of 20m, the gauge pressure is 200,000 Pa. If I go to a depth of 40m, will the gauge pressure double? Will the absolute pressure double?
Answer #1

Absolute pressure = Patm+ syn. P = Patm+ (89x20) P2 = Patm+ (89840) No, absolute pressure wont double. & density g gravitati